Zoron is 10 years older than his sister. In three years, he will be twice as old as she will be then. How old are they now?
Answer by greenestamps(13200)



First a typical formal algebraic setup and solution....

Let x = sister's age
then x+10 = Zoron's age
x+3 = sister's age in three years
(x+10)+3 = x+13 = Zoron's age in three years

Zoron's age in three years will be twice his sister's age then:

x+13 = 2(x+3)
x+13 = 2x+6
7 = x

ANSWERS:
Zoron's age now: x+10 = 17
sister's age now: x = 7

And now a less traditional and less formal solution, using logical reasoning....

The difference in their ages is 10 years, so when he is twice as old as his sister, their ages will be 10 and 20. That will be 3 years from now, so their ages now are 7 and 17.

ANSWERS (again): 7 and 17